(a) How much does a 70.0 kg person weigh on the Earth?
(b) How much would she weigh on the Moon (gmoon = 0.166g)?
(c) How much would she weigh on a neutron star where gstar = 1.43×10[sup]11[/sup]g?
1
Expert's answer
2010-11-23T19:22:51-0500
The weight of an object is W = mg, where g is the gravitational acceleration on the on the planet, so
1. On the Earth the weight is W = 70*10 = 700 N
2. On the Moon W = 70* 0.166*10 = 116.2 N
3. On the neutron star W = 70 * 1.43 *1011*10 =1.001x 1014 N.